OpenStudy (studybird123):

his prism has a volume of 240 cm3. What would the volume of the prism be if each dimension was halved? (Scale factor is 1/2.) A. 60 cm^3 B. 120 cm^3 C. 480 cm^3 D. 30 cm^3

OpenStudy (anonymous):

if each dimension is halved then new volume will be 240*(1/2)^3=240/8 = 30

OpenStudy (studybird123):

Ok thank you!! So muchh!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thx